OpenStudy (anonymous):

solve (1/x^2)+(1/x)=2

OpenStudy (anonymous):

Multiply both sides by \(x^2\): \[x^2\left(\frac{1}{x^2}+\frac{1}{x}\right)=2x^2\\ 1+x=2x^2\\ 2x^2-x-1=0\]

OpenStudy (anonymous):

This can be factored.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw